Fermi National Laboratory


Accelerator Update

The integrated luminosity for last week was 7.62 pb-1.

Monday February 2
The day shift began with Recycler experts conducting studies, with MiniBooNE taking beam, and with the Antiproton Source (Pbar) stacking.

At 8:33 AM, Operations put the Tevatron (TeV) through a dry squeeze and then at 9:15 AM into shot setup.

Operations established store 3214 at 11:28 AM with an initial luminosity of 55.42E30. This set a New Luminosity Record. Pbar resumed stacking.

Switchyard and Main Injector (MI) experts began a study period at 1:31 PM.

Intermittent Booster interlock inhibits began halting beam at 3:54 PM. Operators investigated.

At 4:25 PM, a MI RF station (MIRF13) tripped off for the first of three times within ten minutes. It reset each time.

A MI expert completed her studies at 4:54 PM.

Two Booster RF stations (BRF13 & 14) tripped off at 7:35 PM due to a vacuum burst. They reset okay.

A Switchyard expert completed his studies at 7:52 PM.

Tuesday February 3
The midnight shift began with Operations monitoring store 3214, with Pbar stacking, and with MiniBooNE taking beam.

At 12:40 AM, MiniBooNE tripped off for the first of three times over an hour. The trips were due to a problem with control module for a power supply (E:H866).

Operators noted at 2:11 AM that the MI BPMs weren't responding. Without the BPMS operators could not close the orbit. Reboots did no good at first, but eventually, after several reboots, resolved the problem.

MIRF13 began tripping again at 3:36 AM. It tripped off six times over the next four hours. It reset each time. At 9AM, RF technicians investigated the problem.

A Switchyard expert began a study period at 9 AM.

Operations reported at 10:25 AM that the H- Source had been tripping off. A Linac expert investigated.

At 11:14 AM, a Linac RF station (KRF4) tripped off and wouldn't reset. Operators investigated, but could get the station to stay on. Experts arrived and began repairs.

H- Source trips seemed to be causing 400 MeV area Linac trips at 11:51 AM. After a lot of investigation an expert discovered a problem with a nitrogen valve that was supposed to keep a beam stop open. Techs had this valve repaired around 4:30 PM.

At 2:24 PM, a MI tech began a LCW repair on the power supply for MIRF16. He completed his repairs an hour later.

Operations halted stacking at 2:33 PM so Switchyard experts could conduct a Septa scan. This scan also put a hold on MiniBooNE beam. The scan lasted about an hour.

Recycler experts finished their studies at 5:48 PM.

At 6:35 PM, Switchyard experts sent beam to the Switchyard dump on their first pulse of beam. All Switchyard studies finished at 8 PM.

Operations terminated store 3214 at 9:10 PM and then put the TeV back into shot setup at 9:48 PM.

At 10:06 PM, there was a coalescing problem with injecting the first bunch of protons. This put shot setup on hold while experts investigated. They suggested at 11:19 PM that Operations put the TeV through another dry squeeze for fifteen minutes.

At 11:38 PM, operators and the Duty Electrician investigated a problem with a breaker for two power supplies (M:HV200 & 202) at TeV sector F23 service building.

Operations began a new shot setup at 11:42 PM.

Wednesday February 4
The midnight shift began with the TeV in shot setup.

The Duty Electrician completed his repairs on the F23 breaker at 1:19 AM.

Since many Pbar power supplies in the AP1 line were off due to the F23 breaker repair, operators conducted a Pbar beamline tune-up at 1:34 AM. By 2 AM they were reporting a problem with getting beam into the Accumulator. Operations consulted experts and at 4:42 AM, per their suggestions, manually adjusted kicker settings. This allowed beam tune-up to continue.

At 4:15 AM, Operations established store 3217 with an initial luminosity of 55.50E30. This set yet another New Luminosity Record. Pbar resumed stacking.

Today's Plans
The plan for today is to continue with stacking and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/4/2004   email Fermilab